# Multi-wavelength Study of Dust Emission in the Young Edge-on Protostellar Disk HH 212

### v1

@@ 5.1 Opacity prescription
Figure 8: The relation between dust absorption opacities $\mathrm{\kappa_{\nu,abs}}$ and wavelength $\mathrm{\lambda}$. The black dots are the best-fit value from the PDO model. The dark/light shadows stand for the noise uncertainties from our fitting process (Table 3) and absolute uncertainties contributed by central stellar mass and Toomre parameter. The black-dashed line shows the power-law fitting result from the PDO model using $\mathrm{\kappa_{\nu,abs}}$ between ALMA band 7 and VLA band Q. The green dots and shadows are the estimated value and absolute uncertainties of $\mathrm{\kappa_{\nu,abs}}$ in Paper 1. The orange triangle is $\mathrm{\kappa_{\nu,abs}}$ from Beckwith et al. (1990) with $\mathrm{\beta}$=1. The red/purple solid lines denote the $\mathrm{\kappa_{\nu,abs}}$ of DIANA $\mathrm{a_{max}=50\mu m}$ and DSHARP $\mathrm{a_{max}=130\mu m}$. The red/purple dashed lines are the $\mathrm{\kappa_{\nu,eff}}$ of $\mathrm{a_{max}=75\mu m}$ of DSHARP and DIANA dust model.
@@ 5.2 Albedo prescription
In summary, by comparing the distribution of effective albedos, $\mathrm{\omega_{\nu,eff}}$, under different dust assumptions, we can roughly constrain the upper limit of the dust grain size to $\mathrm{a_{max}=130\mu m}$. Furthermore, the $\mathrm{\omega_{\nu,eff}}$ values of the PDO model closely match those of the $\mathrm{a_{max}=75\mu m}$ DSHARP and DIANA dust models, which is consistent with the $\mathrm{a_{max}=75\mu m}$ inferred from the ALMA Band 7 polarization observations (Lee et al., 2021a). However, the $\mathrm{\kappa_{\nu,abs}}$ values for the $\mathrm{a_{max}=75\mu m}$ DSHARP and DIANA dust models differ significantly from those of the PDO model (Figure 8). This discrepancy suggests that the grain composition and porosity in this disk may differ from those assumed in the DSHARP and DIANA models.
@@ 6 Conclusion
    The best-fit absorption opacities of the PDO model between ALMA band 7 to VLA band Q lie between those from DSHARP $\mathrm{a_{max}=130\mu m}$ and DIANA $\mathrm{a_{max}=50\mu m}$ but are closer to the latter. The opacity spectral index $\mathrm{\beta_{B7-BQ}}$ obtained with a value of $\sim$1.2 is between 1.7 and 1.0, inferring a dust grain size between typical ISM (Draine, 2006) and mm-sized grains (Beckwith et al., 1990).
@@
    This paper makes use of the following ALMA data: ADS/JAO.ALMA# 2012.1.00122.S, 2015.1.00024.S, 2017.1.00044.S, 2017.1.00712.S, and 2019.1.00037.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Y.-C.H. and C.-F.L. acknowledge grants from the National Science and Technology Council of Taiwan (110-2112-M-001-021-MY3 and 112-2112-M-001-039-MY3) and the Academia Sinica (Investigator Award AS-IA-108-M01). Z.Y.D.L. acknowledges support from NASA 80NSSCK1095, the Jefferson Scholars Foundation, the NRAO ALMA Student Observing Support (SOS) SOSPA8-003, the Achievements Rewards for College Scientists (ARCS) Foundation Washington Chapter, the Virginia Space Grant Consortium (VSGC), and UVA research computing (RIVANNA). Z.Y.L. is supported in part by NASA 80NSSC20K0533 and NSF AST- 2307199 and AST-1910106. J.J.T. acknowledges support from NASA XRP 80NSSC22K1159.

### v2

@@ 5.1 Opacity prescription
Refer to caption .
Figure 8: The relation between dust absorption opacities $\mathrm{\kappa_{\nu,abs}}$ and wavelength $\mathrm{\lambda}$. The black dots are the best-fit value from the PDO model. The dark/light shadows stand for the noise uncertainties from our fitting process (Table 3) and absolute uncertainties contributed by central stellar mass and Toomre parameter. The black dash-dot line shows the power-law fitting result from the PDO model using $\mathrm{\kappa_{\nu,abs}}$ between ALMA band 7 and VLA band Q. The green dots and shadows are the estimated value and absolute uncertainties of $\mathrm{\kappa_{\nu,abs}}$ in Paper 1. The orange triangle is $\mathrm{\kappa_{\nu,abs}}$ from Beckwith et al. (1990) with $\mathrm{\beta}$=1. The red/purple solid lines denote the $\mathrm{\kappa_{\nu,abs}}$ of DIANA $\mathrm{a_{max}=50\mu m}$ and DSHARP $\mathrm{a_{max}=130\mu m}$. The red/purple/brown dashed lines are the $\mathrm{\kappa_{\nu,eff}}$ of $\mathrm{a_{max}=75\mu m}$ of DSHARP, DIANA and the mixture of these two dust models, respectively
@@ 5.2 Albedo prescription
In summary, by comparing the distribution of effective albedos, $\mathrm{\omega_{\nu,eff}}$, under different dust assumptions, we can roughly constrain the upper limit of the dust grain size to $\mathrm{a_{max}=130\mu m}$. Furthermore, the $\mathrm{\omega_{\nu,eff}}$ values of the PDO model closely match those of the $\mathrm{a_{max}=75\mu m}$ DSHARP and DIANA dust models, which is consistent with the $\mathrm{a_{max}=75\mu m}$ inferred from the ALMA Band 7 polarization observations (Lee et al., 2021a). However, the $\mathrm{\kappa_{\nu,abs}}$ values for the $\mathrm{a_{max}=75\mu m}$ DSHARP and DIANA dust models differ significantly from those of the PDO model (Figure 8). This discrepancy suggests that the grain composition and porosity in this disk may differ from those assumed in the DSHARP and DIANA models. We also calculate $\mathrm{\kappa_{\nu,abs}}$ and $\mathrm{\omega_{\nu,eff}}$ for dust with $\mathrm{a_{max}=75\mu m}$ assuming a mixture with equal mass fractions of the DSHARP and DIANA models. The resulting values of these quantities are closer to those in the PDO dust model compared to DSHARP and DIANA with $\mathrm{a_{max}=75\mu m}$. Detailed modeling of the new dust composition model and other dust properties for Class 0/I disks are crucial to better describing the PDO dust model.
@@ 6 Conclusion
    The best-fit absorption opacities of the PDO model between ALMA band 7 and VLA band Q lie between those from DSHARP $\mathrm{a_{max}=130\mu m}$ and DIANA $\mathrm{a_{max}=50\mu m}$ but are closer to the latter, indicating a different dust composition in Class 0/I disks. The opacity spectral index $\mathrm{\beta_{B7-BQ}}$ obtained with a value of $\sim$1.2 is between 1.7 and 1.0, inferring a dust grain size between typical ISM (Draine, 2006) and mm-sized grains (Beckwith et al., 1990).
@@
    We thank the anonymous referee for useful comments. This paper makes use of the following ALMA data: ADS/JAO.ALMA# 2012.1.00122.S, 2015.1.00024.S, 2017.1.00044.S, 2017.1.00712.S, and 2019.1.00037.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Y.-C.H. and C.-F.L. acknowledge grants from the National Science and Technology Council of Taiwan (110-2112-M-001-021-MY3 and 112-2112-M-001-039-MY3) and the Academia Sinica (Investigator Award AS-IA-108-M01). Z.Y.D.L. acknowledges support from NASA 80NSSCK1095, the Jefferson Scholars Foundation, the NRAO ALMA Student Observing Support (SOS) SOSPA8-003, the Achievements Rewards for College Scientists (ARCS) Foundation Washington Chapter, the Virginia Space Grant Consortium (VSGC), and UVA research computing (RIVANNA). Z.Y.L. is supported in part by NASA 80NSSC20K0533 and NSF AST- 2307199 and AST-1910106. J.J.T. acknowledges support from NASA XRP 80NSSC22K1159.